# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 4.1 Communication: Face-to-Face
Description of PerpExplore-F2F: Given a ring $R$ with a byzantine black hole node $v_{b}$, a set of agents $A=\{a_{1},a_{2},\dots,a_{k}\}$, where $k=\lceil\log(n-1)\rceil+3$, are initially co-located at a safe node, which we term as $home$. Note that, the agents have no knowledge about the position of $v_{b}$. The agents have the knowledge of $n$ (which is the total number of nodes in $R$), whereas they know that $home$ is safe, as it is the initial position of the set of agents, $A$. Moreover, they have no idea about the remaining $n-1$ nodes of $R$, as $v_{b}$ can be any node among these $n-1$ nodes. We call this arc containing these $n-1$ nodes as the suspicious region, and it is denoted by $S$ with cardinality $|S|$.
The main idea of the PerpExplore-F2F algorithm is from the $home$ only two agents, say $a_{1}$ and $a_{2}$ starts exploring the ring perpetually. The algorithm goes like this, one agent (preferably the agent with lowest Id, among the set of agents which are yet to be consumed by $v_{b}$) moves clockwise until it reaches the ’middle node’ of the suspicious region $S$, while the other agent (preferably the agent with second lowest Id, among the set of agents which have not been consumed by $v_{b}$) also performs the same movement, but along counter-clockwise direction, after which they return to the $home$ and this process perpetually continues. Now while executing this movement by these two agents, there can be two possible scenarios, the moment at least one agent gets destroyed by the byzantine black hole, $v_{b}$. Firstly, both the agents can get destroyed (this happens only when $v_{b}$ is the middle node of $S$), whereas secondly, only one agent is destroyed while the other agent following its usual course of algorithm, returns to $home$. For the initial case, the remaining agents at home can uniquely determines the byzantine black hole, $v_{b}$, as the middle node of $S$ and consecutively starts exploring the ring while avoiding $v_{b}$. For the latter case, the agents at the $home$ along with the agent that has returned to $home$, eventually gets to know that, the other agent has not returned and which in turn concludes that the arc, along which the missing agent was exploring contains the node $v_{b}$. This way our algorithm guarantees that, in the worst case, the size of the suspicious region decreases by a factor of at least half of the previous size, in the expense of one agent. So this implies, when $\lceil\log n\rceil$ agents are destroyed in $v_{b}$, the size of the suspicious region becomes exactly 1 (where this node in the suspicious region is nothing but $v_{b}$). In this case the remaining alive agent can determine the exact position of $v_{b}$ and hence, can perpetually explore the ring $R$ avoiding $v_{b}$.
Next we will give a brief description of our algorithm PerpExplore-F2F. The algorithm works as follows, initially each agent at $home$ starts from state Initial0. In this state, an agent initializes the following variables: $S_{lft}$ to 1, $S_{rgt}$ to $n-1$,where $S_{lft}$ and $S_{rgt}$ denotes the distance to the left and the right end nodes of the suspicious region $S$ from the $home$, in the clockwise direction. Also in this state an agent declares its current node as the $home$. Note that, the agents stay at this state only for one round at the beginning of the execution of our algorithm, after which they enter the state Initial1. Now, after an agent enters the state Initial1, it calculates $|S|$ (where $|S|=|S_{rgt}-S_{lft}+1|$), initializes $Tnodes$ to 0 (where $Tnodes$ is the distance an agent traverses in order to reach its destination node). Next, it checks the variables $|S|$ and $k$ (where $k$ is the number of agents present at the $Current-Node$, i.e., $home$). If it finds $|S|=2$ and $k=2$, i.e., the length of the suspicious region is 2 and the total number of agents present at $home$ is also 2, then both the agents execute the sequence $SEQ0$ of states, where $SEQ0$ is defined as follows:
[…]
According to this sequence, the agent moves in to state Detection (after satisfying both the conditions $|S|=2$ and $k=2$ in state Initial1). In this state, the lowest Id agent is instructed to move clockwise and reach its destination position $S_{lft}$, whereas, the other agent reaches its destination $S_{rgt}$ along a counter-clockwise direction, while after each movement updating $Tnodes=Tnodes+1$. Now, after reaching their respective positions, the agents return back to $home$ along the same path in which they reached their destination nodes, again incrementing $Tnodes$ by 1 after each move. So, the total number of rounds required by the agent, starting from Initial1 in order to reach $home$ is $Tnodes+1$. After reaching $home$, they are further instructed to wait for $22n-Tnodes+$ 2 rounds. In the following round it again checks whether, #$\#Agent$ at $home$ is still $k$ i.e., 2, if so then again it changes to state Initial1, and the sequence $SEQ0$ is executed repeatedly. Otherwise, if the current agent is of lowest Id then declares the byzantine black hole node is at the location of $S_{rgt}$, otherwise if the current agent is not of lowest Id, then it declares that this byzantine black hole node is at $S_{lft}$. Irrespective of which, they continue performing perpetual exploration of $R$, by avoiding this black hole node. Hence, the total number of rounds required to complete the execution of this sequence is $42n+$ 4.
On the contrary, if either $|S|\neq 2$ or $k\neq 2$. Note that, if $|S|=1$, then this node is nothing but the byzantine black hole node, and the remaining agents (such agent always exists, refer Lemma 4.11) detect this, and further continue perpetual exploration, avoiding this node. Otherwise, for $|S|\geq 2$ and $k\geq 3$ in state Initial1, the agents update the following variables: $k$ to $k-2$ (where $k-2$ is the number of agent which enters the next state Stagnant), locates the next destination node, where it calculates the variable $Dest=S_{lft}+\lceil\frac{|S|}{2}\rceil-1$, which is the clockwise distance of the destination node from $home$, $W_{time}$ to 00 (where $W_{time}$ signifies the waiting time at destination node). Agents in Initial1 state can change their state in to one of the following three states: Forward-Left, Forward-Right and Stagnant. An agent moves in to Forward-Left, if it is the agent with lowest Id among all the agents present at $home$, whereas an agent moves in to Forward-Right, if the agent is having second lowest Id among all the agents present at $home$. Lastly, an agent enters the state Stagnant, if it is neither with lowest Id nor with second lowest Id.
According to the state an agent enters after Initial1, we define three sequence of states $SEQ1$, $SEQ2$ and $SEQ3$, respectively, which an agent can follow in the exact order after Initial1 state, based on whether the agent enters Forward-Left or Forward-Right or Stagnant. If the agent enters the state Forward-Left then it starts following $SEQ1$, or if it enters the state Forward-Right then it follows $SEQ2$ or if it enters Stagnant state then it follows $SEQ3$, respectively. Note that, in each sequence of states the agent eventually reaches back to Initial1, if not it is destroyed at the node $v_{b}$, while performing any of the states: Forward-Left (resp, Forward-Right) or Wait-Left (resp, Wait-Right) or Backtrack-Left (resp, Backtrack-Right). Also note that if one agent stated executing sequence $SEQ1$, then there must be another agent that executes sequence $SEQ2$ and the other agents execute the sequence $SEQ3$ starting from the same round (refer Line 10-17 of Algorithm  1). We define $SEQ1$, $SEQ2$ and $SEQ3$ as follows:
[…]
Our algorithm guarantees that, each agent starting from Initial1 again enters the state Initial1 at the same round, after following any of these three sequences.
Definition 4.1 (Iteration).
We define an iteration of our algorithm as the number of rounds required to execute any one of these sequences (i.e., $SEQ0$, $SEQ1$, $SEQ2$ and $SEQ3$) by an agent.
Note that, each iteration has exactly $42n+$ 4 rounds. Each agent performs these iterations until this byzantine black hole is detected, after which, they explore the ring $R$ perpetually just by avoiding this black hole node.
Next, we describe how the agents execute the sequences in order to perform the PerpExplore-F2F algorithm in one iteration.
1 Input: $n,k$;
2 States: {Initial0, Initial1, Stagnant, Forward-Left, Forward-Right, Backtrack-Left, Backtrack-Right, Wait-Left, Wait-Right, Wait1-Left, Wait1-Right}
3
4In State Initial0:
5 $S_{lft}=$ 1, $S_{rgt}=$ $n-1$;
6
7Declare $Current-Node$ as $home$ and update to state Initial1.
8
9In State Initial1:
10 $|S|=|S_{rgt}-S_{lft}+1|$, $Tnodes=0$;
11 if $S_{lft}\neq S_{rgt}$ then
12       if $|S|=2\wedge k=2$ then
13             Move to state Detection.
14
15      else
16             $k=k-2$, $Dest=S_{lft}+\lceil\frac{|S|}{2}\rceil-1$, $W_{time}=0$;
17             if $LowestID$ then
18                   Update state to Forward-Left.;
19
20             else if $SecondLowestID$ then
21                   Update to state Forward-Right.;
22
23             else
24                   Move to state Stagnant.;
25
26
27
28else
29       Declare $S_{lft}$ as the byzantine black hole location and continue exploration avoiding that node.
30
31
32In State Stagnant:
33 Wait at the $Current-Node$ for $22n+$ 2 rounds.
34 if #$\#Agent=k$ then
35       Declare $Dest$ is the byzantine black hole location and continue exploration avoiding that node.
36
37 else if #$\#Agent=k+1$ then
38       Update $k=k+1$ and collect the updated $S_{lft}$, $S_{rgt}$ and update accordingly.
39       Move to state Initial1.
40
41 else
42       Update $k=k+2$ and move to state Initial1.
43
44
[…]
47
48In State Wait-Left:
49
50if $W_{time}<n-Tnodes$  then
51       $W_{time}++$
52else
53       Enter state Backtrack-Left and move $right$
54
55
56In State Wait-Right:
57
58if  $W_{time}<n-Tnodes$  then
59       $W_{time}++$
60else
61       Enter state Backtrack-Right and move $left$
62
63
Algorithm 1 PerpExplore-F2F
Execution of the sequence $SEQ1$: As discussed earlier, if the agent is of lowest Id in the in the state Initial1, then it is instructed to change their state to Forward-Left. In state Forward-Left, an agent reaches the destination node, which is $Dest$ distance away from $home$, along the clockwise direction, while after each move incrementing the variable $Tnodes$ by 1. Now, the moment it reaches the destination node, it is further instructed to change its state to Wait-Left. Note that, an agent stays in Forward-Left state for exactly $Tnodes+1$ rounds while following the sequence $SEQ1$ in an iteration. After changing its state to Wait-Left, the agent waits at the destination node for $n-Tnodes-1$ rounds, after which it is again instructed to move in to another state, which is Backtrack-Left while already starting to move one hop along counter-clockwise direction. Observe that an agent spends exactly $n-Tnodes$ round in state Wait-Left while executing $SEQ1$ in an iteration. Next in the state Backtrack-Left, an agent immediately starts moving along counter-clockwise direction, until the agent reaches the $home$ node, which is $Tnodes-1$ distance apart from its current node. While it reaches $home$, it again changes to state Wait1-Left. An agent stays at the Backtrack-Left state for exactly $Tnodes$ rounds. State Wait1-Left, instructs the agent to wait at the current node, i.e., $home$ for $n-Tnodes$ many rounds. This waiting time is enough for another agent that started executing sequence $SEQ2$ to return back to the $home$ if not destroyed by the byzantine black hole. After waiting, the agent checks the number of agent present at the current node. If #$\#Agent$ increases to $k+1$ (where $k$ agents are in state Stagnant), then in that case, it updates $k$ to $k+1$ and understands that the second lowest Id agent has not returned to $home$. In this situation, the only possibility of the other agent not returning back is that it has entered the black hole, along its traversal. In this situation it updates $S_{lft}$ to the destination node, which is $Dest$ distance apart from $home$ (i.e., basically the suspicious region becomes half, because initially $S_{lft}$ was 1, whereas now it is updated to $Dest$), and also communicates this information to the remaining $k-1$ agents at $home$, and directly moves into state Initial1.
42
[…]
45 In State Wait1-Left:
46 Update $d=0$ and wait at the $Current-Node$ for $n-Tnodes$ rounds.
47 if #$\#Agent=k+1$ then
48       Update $k=k+1$ and $S_{lft}=Dest$ and communicate $S_{lft}$ to remaining agents at $Current-Node$.
49       In the next round move to state Initial1.
50
51else
52       Update $k=k+2$.
53       In the next round move to state Initial1.
54
55
56In State Wait1-Right:
57 Update $d=0$ and wait at the $Current-Node$ for $n-Tnodes$ rounds.
58 if #$\#Agent=k+1$ then
59       Update $k=k+1$ and $S_{rgt}=Dest$ and communicate $S_{rgt}$ to remaining agents at $Current-Node$.
60       In the next round move to state Initial1.
61
62else
63       Update $k=k+2$.
64       In the next round move to state Initial1.
65
66
67In State Detection:
68 If lowest Id then move to $S_{lft}$ clockwise, otherwise move to $S_{rgt}$ counter-clockwise, and after each move, update $Tnodes=Tnodes+1$.
69 Return back to $home$ and again update $Tnodes=Tnodes+1$ after each move.
70 if $Current-Node$ is $home$ then
71       Wait at $home$ for $22n-Tnodes+$ 2 rounds.
72       if #$\#Agent=k$ then
73             Move to state Initial1.
74
75      else
76             If lowest Id then declare byzantine black hole is $S_{rgt}$, otherwise declare byzantine black hole is $S_{lft}$, and then continue exploration avoiding this node.
77
78
79
On the other hand, if #$\#Agent$ at $home$ increases from $k$ to $k+2$ (i.e., both lowest Id and second lowest Id has returned back to $home$) then the agent simply waits for one round, after which it directly moves to state Initial1. Here note that an agent stays in the state Wait1-Left for $n-Tnodes+2$ rounds. Hence for an agent executing sequence $SEQ1$, the length of an iteration is $42n+$ 4.
Execution of the sequence $SEQ2$. This execution if followed by the agent, which in state Initial1 has the second lowest Id. In this sequence, the agent from Initial1 moves to state Forward-Right. The state Forward-Right is exactly the same as the state Forward-Left which has been discussed above, the only difference is that the agent moves along a counter-clockwise direction, instead of clockwise direction. After this state, the agent moves to state Wait-Right, in this state again similar to the earlier state Wait-Left, the agent waits for $n-Tnodes-1$ many rounds at the destination node. After which it directly moves along clockwise direction for one hop and moves to state Backtrack-Right. This state resembles Backtrack-Left, instead the fact that in this state the agent moves along clockwise direction for $Tnodes-1$ distance, in which case it reaches $home$. After reaching $home$, it again waits for $n-Tnodes$ rounds while in state Wait1-Right, and after which performs similar checks as discussed earlier in Wait1-Left, but here, if after the waiting period, the total number of agents at $home$ is $k+1$ (where $k$ many agents are in state Stagnant), it similarly modifies the suspicious region by updating $S_{rgt}$ to $Dest$ (while earlier $S_{lft}$ was updated to $Dest$ in state Wait1-Left). Lastly, it ends this sequence $SEQ2$ by moving in to state Initial1 again. Similar to the length of an iteration for an agent executing $SEQ1$, the length of an iteration for any agent executing $SEQ2$ is again $42n+$ 4.
Execution of the sequence $SEQ3$. This execution is followed by the agent, which in state Initial1 finds that it is neither of lowest Id nor of second lowest Id. In this state, the agent is instructed to move to state Stagnant. While in state Stagnant, it waits for $22n+$ 2 rounds at $home$, after which in $(2n+3)$-th round it checks the number of agent present at $home$. If ##\##Agent is $k$ (i.e., only the agents which moved to state Stagnant), then the agent understands that the lowest Id agent (which started sequence $SEQ1$) and second lowest Id agent (which started sequence $SEQ2$), both has entered the black hole, and since $R$ has only one black hole node, so this scenario of both agents entering the black hole can only occur when the node is common to both of these agents. This common node is nothing but the destination node, hence if this scenario happens, then the agent in state Stagnant, concludes that the black hole is none other than the destination node, and continues perpetual exploration avoiding this node. On the contrary, if #$\#Agent$ is $k+1$ (i.e., one among the agents executing $SEQ1$ and $SEQ2$ has returned, whereas the other agent has entered the black hole). In this scenario, updates either $S_{lft}$ or $S_{rgt}$ (based on the fact, whether the agent executing $SEQ1$ or the agent executing $SEQ2$ has returned), and then moves to state Initial1 while updating $k$ to $k+1$, ending sequence $SEQ3$. On the other hand, if both the agents executing $SEQ1$ and $SEQ2$ return to $home$ (i.e., ##\##Agent is $k+2$), then the agents in state Stagnant simply moves to state Initial1 while updating $k$ to $k+2$, and ending the sequence $SEQ3$. Note that here also, the iteration is of length $42n+$ 4 rounds, which leads to the following observation.
{observation}
For any $i$ ($i>0$), the $i$-th iteration for each of the $k$ alive agents, starts at the same round and at the same node, i.e., $home$.
{observation}
For any $i$ ($i>0$), each agent moves to the state Initial1 at the end of this $i$-th iteration.
{observation}
If $|S|=2$ and $k=2$, then both these agents follow the sequence $SEQ0$, whereas, if $|S|\geq 2$ and $k\geq 3$, then one agent follows $SEQ1$, another agent follows $SEQ2$, whereas the remaining agents follow the sequence $SEQ3$. That is,during an iteration if an agent follows $SEQ0$ then no agent follows $SEQ1$, $SEQ2$ or $SEQ3$ in that iteration, and vice-versa.
4.1.1 Correctness and Complexity
In this section we give the correctness and complexity analysis of our algorithm PerpExplore-F2F.
Lemma 4.2.
During any iteration, if exactly one agent gets destroyed by the byzantine black hole, then at the end of this iteration, the length of the suspicious region becomes at most $\lceil\frac{|S|}{2}\rceil+1$, where $|S|\geq 2$ is the length of the suspicious region at the start of this iteration.
Proof 4.3.
Let us consider at the $i$-th iteration of our Algorithm 1, the suspicious region be $S$ (where, $|S|>1$), so the agent starts from the state Initial1, and can encounter one of the following scenarios: first, $|S|=2$ (i.e., $S=\{S_{lft},S_{rgt}\}$) and $k=2$, second, $|S|\geq 2$ and $k>2$ (refer line 8 of Algorithm 1).
If the first scenario is satisfied, then both the agents execute the sequence $SEQ0$ (refer description of PerpExplore-F2F), in which they perpetually visit the nodes $S_{lft}$ and $S_{rgt}$, respectively and return back to $home$, until one of these agent fails to return $home$, i.e., enters the byzantine black hole (refer state Detection in Algorithm 1). So, while the other agent returns $home$ (which was visiting $S_{rgt}$, without loss of generality), it is instructed to wait for $22n-Tnodes+$ 2 rounds (refer line 63 of Algorithm 1). After waiting, it finds that the agent visiting $S_{lft}$ is yet to return $home$, then this phenomenon helps the alive agent conclude that the agent visiting $S_{lft}$ must have entered the black hole (because $22n-Tnodes+$ 2 rounds is sufficient number of rounds for any agent visiting the node $S_{lft}$ to return back to $home$, if not consumed by the black hole), which is none other than the node $S_{lft}$ (as it is the only node in the suspicious region that this agent visits). Hence, in this case the black hole node is detected and also $|S|$ becomes 1 at the end of $i$-th iteration, which satisfies our claim.
On the otherhand, if $|S|\geq 2$ and $k>3$, the agents are either following $SEQ1$ or $SEQ2$ or $SEQ3$, among them exactly one agent follows $SEQ1$ (say, $a_{1}$), exactly one agent follows $SEQ2$ (say, $a_{2}$), whereas the rest of the agent follows $SEQ3$ (where these agents stay at $home$ during this iteration, refer the description of execution of $SEQ3$). This means that the agent which can be destroyed by the black hole can be either $a_{1}$ or $a_{2}$ (as $home$ is always a safe node). If $a_{1}$ is destroyed, then $a_{2}$ during this iteration, after returning back to $home$ (while in state Backtrack-Right), waits at $home$ for $n-Tnodes$ many rounds (while in state Wait1-Right) for $a_{1}$ to return. Since, $a_{1}$ has already been consumed by the black hole, so it fails to return to $home$ within the waiting time of $a_{2}$. This observation of $a_{2}$ at $home$ instigates it to modify the suspicious region by updating $S_{rgt}$ to $Dest$, and $a_{2}$ communicates this updated information to the rest of the agents executing $SEQ3$ at $home$. It is because, $a_{1}$ only explores the nodes of the suspicious region that lies between $S_{lft}$ and $Dest$. Let $S_{new}$ be the modified suspicious region, then $|S_{new}|=|Dest-S_{lft}+1|=|S_{lft}+\lceil\frac{|S|}{2}\rceil-1-S_{lft}+1|=\lceil\frac{|S|}{2}\rceil<\lceil\frac{|S|}{2}\rceil+1$.
Now consider the other case, where instead of $a_{1}$, $a_{2}$ be the agent which enters the black hole during an iteration. In this situation as well, similar to the argument explained above in case of $a_{2}$, $a_{1}$ will get to know $a_{2}$ has entered the black hole during the iteration, while $a_{1}$ waits for $n-Tnodes$ rounds at $home$. Observe, $a_{2}$ traverses the nodes of the suspicious region that lies between $Dest$ and $S_{rgt}$. So, $a_{1}$ updates $S_{lft}$ to $Dest$ in order to get the updated suspicious region $S_{new}$ and communicate with the rest of the agents executing $SEQ3$ at $home$. This implies $|S_{new}|=|S_{rgt}-Dest+1|=|S_{rgt}-S_{lft}-\lceil\frac{|S|}{2}\rceil+2|=||S|-\lceil\frac{|S|}{2}\rceil+1|\leq\lceil\frac{|S|}{2}\rceil+1$.
So, in each scenario, our claim holds.
{observation}
Note that, if $|S|$ is odd then $|S_{new}|=\lceil\frac{|S|}{2}\rceil$, otherwise, $\lceil\frac{|S|}{2}\rceil\leq|S_{new}|\leq\lceil\frac{|S|}{2}\rceil+1$.
Lemma 4.4.
If during an iteration, exactly two agents gets destroyed by the byzantine black hole, then in that iteration, the position of the black hole is detected by the remaining alive agents.
Proof 4.5.
This situation of both agents being consumed by the black hole cannot arise when $|S|=2$ and $k=2$, because in this case both agents explore disjoint nodes of the suspicious region, in which only one node can be the black hole.
On the other hand when $|S|\geq 2$ and $k>3$, this situation of two agents entering the black hole, can only arise for the agents $a_{1}$ and $a_{2}$, where $a_{1}$ and $a_{2}$ are following $SEQ1$ and $SEQ2$, respectively. Now, since the node (say $v_{b}$) at distance $Dest$ in the clockwise direction from $home$ is the only common node that both $a_{1}$ and $a_{2}$ visits while executing an iteration. So, if both these agents gets destroyed, then $v_{b}$ must be the node, in which the black hole is located. Now, as per our algorithm, the agents executing $SEQ3$ waits at $home$ for $22n+$ 2 rounds, after which they find that both $a_{1}$ and $a_{2}$ are yet to return $home$, whereas $22n+$ 2 rounds are sufficient for these two agents to return back to $home$ while executing their respective sequences, if they have not entered the black hole. This scenario helps the remaining agent conclude that both $a_{1}$ and $a_{2}$ has entered the black hole, and the node $v_{b}$ is the exact position of the byzantine black hole.
Lemma 4.6.
Let $|S_{0}|$ be the initial length of the suspicious region where $|S_{0}|=2^{p}+l$, $p\geq 1$ and $0\leq l<2^{p}$. Let $S_{i}$ be the suspicious region after it is modified for the $i-$th time. Then, $|S_{i}|\leq\lceil\frac{|S_{0}|}{2^{i}}\rceil+2$.
Proof 4.7.
Let $|S_{0}|$ be the initial length of the suspicious region where $|S_{0}|=2^{p}+l$, $p\geq 1$ and $0\leq l<2^{p}$. Let $S_{i}$ be the suspicious region after $i$ modifications. We now prove this lemma by induction. For the base case let $i=1$. Then from Lemma 4.2, $|S_{1}|\leq\lceil\frac{|S_{0}|}{2}\rceil+1<\lceil\frac{|S_{0}|}{2}\rceil+2$. So the lemma is true for $i=1$. Let the lemma be true for $i=j-1$. That is, $|S_{j-1}|\leq\lceil\frac{|S_{0}|}{2^{j-1}}\rceil+2$. Now to prove the lemma we have to show that $|S_{j}|\leq\lceil\frac{|S_{0}|}{2^{j}}\rceil+2$. Now by Lemma 4.2, $|S_{j}|\leq\lceil\frac{|S_{j-1}|}{2}\rceil+1\leq\lceil\frac{\lceil\frac{|S_{0}|}{2^{j-1}}\rceil+2}{2}\rceil+1\leq\lceil\frac{|S_{0}|}{2^{j}}\rceil+2$. Hence, by mathematical induction we can conclude that $|S_{i}|\leq\lceil\frac{|S_{0}|}{2^{i}}\rceil+2$, for any $i\geq 1$
Lemma 4.8.
If $|S_{t}|>2$ , then there exists at least one agent which is following the sequence $SEQ3$ in our Algorithm 1 in any iteration until the next modification of $S_{t}$, where $S_{t}$ is the suspicious region after $t$ modifications .
Proof 4.9.
Let $S_{0}$ be the initial suspicious region. Then, $|S_{0}|=2^{p}+l$ where $p\geq 1$ and $0\leq l<2^{p}$. Let $S_{i}$ be the suspicious region after $i$ modifications of it. Let for some $t>0$, $|S_{t}|>2$. This implies that during all the previous $t$ iterations where the suspicious region is modified, only one agent is consumed by the byzantine black hole. As otherwise $|S_{j}|=1$ for some $0<j<t$, and already the byzantine black hole is detected. This implies after the $t$-th modification, number of agents consumed is exactly $t$. So the remaining number of agent after $t$-th modification of suspicious region is exactly $\lceil\log|S_{0}|\rceil+3-t$ (since $\lceil\log|S_{0}|\rceil+3$, are the total number of agents executing PerpExplore-F2F along $R$). Now in any iteration until the next modification of $S_{t}$, there is an agent that follows $SEQ3$ if $\lceil\log|S_{0}|\rceil+3-t\geq 3\implies t\leq\lceil\log|S_{0}|\rceil$.
Now we suppose, let $t>\lceil\log|S_{0}|\rceil$. Now there are two cases.
Case-I : $|S_{0}|=2^{p}$ (i.e., $l=0$). Then, $t>p$. Now by Lemma 4.6, $|S_{p}|\leq 3$. Also, number of alive agent is exactly 3 after the $p-$th modification. Now if $|S_{p}|$ is not less than or equal to 2, then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p}$. Now , $|S_{p+1}|\leq 2$ ( Observation 4.1.1). Now this is a contradiction as $2<|S_{t}|\leq|S_{p+1}|\leq 2$.
Case-II : Now let us consider the case where $|S_{0}|=2^{p}+l$ where $02^{p}>l>$ 0. Then $t>p+1$. Now by Lemma 4.6, $|S_{p}|\leq 4$. Also, number of alive agent is exactly 4 after the $p-$th modification. Now if $|S_{p}|$ is not less than or equal to 2, then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p}$. Now, $|S_{p+1}|\leq 3$ (by Observation 4.1.1). Again if $|S_{P+1}|$ is $\geq 2$ then the agents will again follow the sequences $SEQ1,SEQ2$ and $SEQ3$ until the next modification of $S_{p+1}$. Now again by Observation 4.1.1, $|S_{P+2}|\leq 2$. Now this again is a contradiction as $2<|S_{t}|\leq|S_{p+2}|\leq 2$.
Hence $t\leq\lceil\log|S_{0}|\rceil$. Thus, there must exist an agent following $SEQ3$ in any iteration after the $t-$th modification of suspicious region until the next one where $|S_{t}|>2$
From the above proof we can have the following corollary
Corollary 4.10.
If $|S_{t}|>2$ where $S_{t}$ is the suspicious region after $t$ modifications, then there is atleast 3 alive agents for all iterations after the $t$-th modification of the suspicious region and until the next modification.
Lemma 4.11.
[…]

### v2

@@ 4.1 Pebble Model of Communication
Brief Description of the Algorithm: Given a ring $R$ with a Byzantine black hole node $v_{b}$, a set of agents, $A=\{a_{1},a_{2},a_{3}\}$ (where we assume, ID of $a_{1}\leavevmode\nobreak\ <$ ID of $a_{2}\leavevmode\nobreak\ <$ ID of $a_{3}$), are initially co-located at $home$. The agents have no knowledge about the position of the node $v_{b}$, the only set of knowledge the agents have are: the total number of nodes in the underlying topology $R$ (i.e., $n$), and also the knowledge that the initial node, i.e., $home$ is safe. So, the remaining arc of $n-1$ nodes in $R$ is a suspicious region (which we term as $S$, where the cardinality of $S$ i.e., $|S|$ denotes the length or size of the suspicious region) for each of the three agents.
The main idea of this algorithm is as follows: initially two agents $a_{1}$ and $a_{2}$ explores the ring perpetually, while $a_{3}$ waits at the $home$. $a_{1}$ and $a_{2}$ explores the ring $R$, executing the rule described as follows. If all the three agents are at $home$ along with two pebbles, then in the $r-$th round ($r\geq 0$), $a_{1}$ moves clockwise with one pebble and it waits at round $r+1$ for $a_{2}$ to arrive. In round $(r+1)$, $a_{2}$, leaving the remaining pebble at $home$ moves to the next node along clockwise direction to meet $a_{1}$. In the subsequent round, i.e., $(r+2)-$th round, $a_{1}$ after meeting $a_{2}$ moves again to the next node in the clockwise direction and waits for 3 rounds (i.e., till $(r+5)-$th round) for $a_{2}$ if and only if $a_{1}$ is not at $home$. In the mean time at $(r+3)-$th round, $a_{2}$ leaves its current node, moves one step in counter-clockwise direction, collects the left behind pebble. Then in $(r+4)-$th round $a_{2}$ moves again to the earlier node in clockwise direction. Subsequently, in $(r+5)-$th round, $a_{2}$ again leaves the pebble at its current node, and moves in a clockwise direction, to meet $a_{1}$ (which is waiting at the current node for $a_{2}$ to arrive), for better reference see Fig. 1. Note that, when $a_{2}$ meets $a_{1}$ outside $home$, the pebble it was carrying is left by $a_{2}$ at the nearest counter-clockwise node of the meeting node (i.e., where $a_{2}$ meets with $a_{1}$). Now from round $(r+6)$ onwards to $(r+9)-$th round, the same execution repeats, as it has occurred between round $(r+2)$ to round $(r+5)$. The only difference is that, if both agents (i.e., $a_{1}$ and $a_{2}$) at round $r+2$ were at a node $v_{1}$, at round $r+6$, they are on the nearest clockwise node of $v_{1}$ (i.e., at $v_{2}$, say). This process continues until $a_{1}$ reaches $home$. In this case $a_{1}$ does not move clockwise, until $a_{2}$ meets it with the pebble it was carrying, in that case, it waits further until $a_{2}$ brings that pebble back at $home$. For this to happen after reaching $home$, 5 rounds are sufficient for waiting. When $a_{2}$ reaches $home$ with the pebble, then all 3 agents are again at the $home$ with 2 pebbles. This way the 3 agents explore the ring. The exploration can hamper if either $a_{1}$ or, $a_{2}$, or a pebble is destroyed by the Byzantine black hole $v_{b}$ while the above mentioned procedure is executed by them.
Figure 1: An execution of PerpExplore-Coloc-Pbl, starting from the configuration where $a_{0}$ and $a_{1}$ are together on a vertex, to the configuration where $a_{0}$ and $a_{1}$ are on the same vertex again, which is the clockwise neighbour of the earlier vertex. The red node is the byzantine black hole, the green node is the $home$ and red boxes are pebbles.
Firstly, let only $a_{1}$ is destroyed at some round $t$. Then there must exists a round $t^{\prime}<t$, when $a_{2}$ was with $a_{1}$ at some node $v_{0}$ (say) for the last time. This means at round $t^{\prime}$, $a_{1}$ moved clockwise to the next node $v_{1}$ along with the pebble it was carrying, and it must have been destroyed before or at the round when $a_{2}$ reaches $v_{1}$ again (in this case $v_{1}$ is $v_{b}$). Now when $a_{2}$ reaches $v_{1}$ there can be two cases. Either $a_{2}$ remains alive or (as the adversary may choose not to destroy $a_{2}$), it can get destroyed as well. For the initial case, $a_{2}$ identifies the Byzantine black hole by not seeing $a_{1}$ there and then it can leave that node and can explore the ring perpetually avoiding the node. The latter case is equivalent to both $a_{1}$ and $a_{2}$ are destroyed by the Byzantine black hole $v_{b}=v_{1}$. Note that since outside $home$ whenever $a_{2}$ reaches the same node of $a_{1}$, it leaves behind a pebble at the nearest counter-clockwise node (here $v_{0}$). For this case, another agent (i.e., $a_{3}$) which was waiting at $home$, finds none of the two exploring agents returns at $home$, even after waiting for a sufficient number of rounds. This incident triggers $a_{3}$ to move clockwise until it finds a pebble (one that is left behind by $a_{2}$ at $v_{0}$). Whenever the pebble is found, $a_{3}$ knows that the next clockwise node is the Byzantine black hole and it starts exploring $R$ avoiding that node.
[…]